# Towards Cross-Tokenizer Distillation: the Universal Logit Distillation Loss for LLMs

### v2

@@ A Legend of Confucius
When confucius came to the earth, the kilin, that strange beast which is the prince of all four-footed animals, and only appears when there is a great man on earth, sought the child and spat out a jade whereon was written: ’son of the watercrystal you are destined to become an uncrowned king!’ and confucius grew up, studied diligently, learned wisdom and came to be a saint. He did much good on earth, and ever since his death has been reverenced as the greatest of teachers and masters. He had foreknowledge of many things and even after he had died, he gave evidence of this.
Why was confucius’s death reverenced as the greatest of teachers and masters?
He did much good on earth.
Naughty Boy
@@
’Oh, let me in! Let me in! I’m cold, and I’m so wet!’ Exclaimed suddenly a child that stood crying at the door and knocking for admittance, while the rain poured down, and the wind made all the windows rattle. ’Poor thing!’ said the old poet, as he went to open the door. there stood a little boy, quite naked, and the water ran down from his long golden hair. He trembled with cold, and had he not come into a warm room he would most certainly have perished in the frightful tempest.
Why did the boy ask to come inside?
He was cold and wet.
@@
Injury severity score (ISS), Glasgow coma score (GCS), and revised trauma score (RTS) are the most frequently used methods to evaluate the severity of injury in blunt trauma patients. ISS is too complicated to assess easily and GCS and RTS are easy to assess but somewhat subjective. White blood cell count (WBC) is an easy, quick and objective test. This study was performed to evaluate the significance of the WBC count at presentation in the blunt trauma patients. 713 blunt trauma patients, who were admitted to the Uludag University Medical Center Emergency Department between 01.04.2000-31.12.2000, were retrospectively evaluated in terms of ISS, GCS, RTS and white blood cell count at presentation. Statistical analysis revealed that WBC was correlated positively with ISS, but negatively with GCS and RTS.
Does the leukocyte count correlate with the severity of injury
The leukocyte count at presentation can be used as an adjunct in the evaluation of the severity of injury in blunt trauma patients.
@@
The aim of this study was to assess the diagnostic value of articular sounds, standardized clinical examination, and standardized articular ultrasound in the detection of internal derangements of the temporomandibular joint. Forty patients and 20 asymptomatic volunteers underwent a standardized interview, physical examination, and static and dynamic articular ultrasound. Sensitivity, specificity, and predictive values were calculated using magnetic resonance as the reference test. A total of 120 temporomandibular joints were examined. Based on our findings, the presence of articular sounds and physical signs are often insufficient to detect disk displacement. Imaging by static and dynamic high-resolution ultrasound demonstrates considerably lower sensitivity when compared with magnetic resonance. Some of the technical difficulties resulted from a limited access because of the presence of surrounding bone structures.
Internal derangement of the temporomandibular joint: is there still a place for ultrasound?
The present study does not support the recommendation of ultrasound as a conclusive diagnostic tool for internal derangements of the temporomandibular joint.
@@
Figures from the British Defence Dental Services reveal that serving personnel in the British Army have a persistently lower level of dental fitness than those in the Royal Navy or the Royal Air Force. No research had been undertaken to ascertain if this reflects the oral health of recruits joining each Service. This study aimed to pilot a process for collecting dental and sociodemographic data from new recruits to each Service and examine the null hypothesis that no differences in dental health existed. Diagnostic criteria were developed, a sample size calculated and data collected at the initial training establishments of each Service. Data for 432 participants were entered into the analysis. Recruits in the Army sample had a significantly greater prevalence of dental decay and greater treatment resource need than either of the other two Services. Army recruits had a mean number of 2.59 (2.08, 3.09) decayed teeth per recruit, compared to 1.93 (1.49, 2.39 p<0.01) in Royal Navy recruits and 1.26 (0.98, 1.53 p<0.001) in Royal Air Force recruits. Among Army recruits 62.7% were from the two most deprived quintiles of the Index of Multiple Deprivation compared to 42.5% of Royal Naval recruits and 36.6% of Royal Air Force recruits.
Is there a differential in the dental health of new recruits to the British Armed Forces?
A significant difference in dental health between recruits to each Service does exist and is a likely to be a reflection of the sociodemographic background from which they are drawn.
@@
11 Appendix - Training Information

### v3

@@ E.2 Few-Shot examples
A Legend of Confucius When confucius came to the earth, the kilin, that strange beast which is the prince of all four-footed animals, and only appears when there is a great man on earth, sought the child and spat out a jade whereon was written: ’son of the watercrystal you are destined to become an uncrowned king!’ and confucius grew up, studied diligently, learned wisdom and came to be a saint. He did much good on earth, and ever since his death has been reverenced as the greatest of teachers and masters. He had foreknowledge of many things and even after he had died, he gave evidence of this. Why was confucius’s death reverenced as the greatest of teachers and masters? He did much good on earth.
@@
Naughty Boy ’Oh, let me in! Let me in! I’m cold, and I’m so wet!’ Exclaimed suddenly a child that stood crying at the door and knocking for admittance, while the rain poured down, and the wind made all the windows rattle. ’Poor thing!’ said the old poet, as he went to open the door. there stood a little boy, quite naked, and the water ran down from his long golden hair. He trembled with cold, and had he not come into a warm room he would most certainly have perished in the frightful tempest. Why did the boy ask to come inside? He was cold and wet.
@@
Injury severity score (ISS), Glasgow coma score (GCS), and revised trauma score (RTS) are the most frequently used methods to evaluate the severity of injury in blunt trauma patients. ISS is too complicated to assess easily and GCS and RTS are easy to assess but somewhat subjective. White blood cell count (WBC) is an easy, quick and objective test. This study was performed to evaluate the significance of the WBC count at presentation in the blunt trauma patients. 713 blunt trauma patients, who were admitted to the Uludag University Medical Center Emergency Department between 01.04.2000-31.12.2000, were retrospectively evaluated in terms of ISS, GCS, RTS and white blood cell count at presentation. Statistical analysis revealed that WBC was correlated positively with ISS, but negatively with GCS and RTS. Does the leukocyte count correlate with the severity of injury The leukocyte count at presentation can be used as an adjunct in the evaluation of the severity of injury in blunt trauma patients.
@@
The aim of this study was to assess the diagnostic value of articular sounds, standardized clinical examination, and standardized articular ultrasound in the detection of internal derangements of the temporomandibular joint. Forty patients and 20 asymptomatic volunteers underwent a standardized interview, physical examination, and static and dynamic articular ultrasound. Sensitivity, specificity, and predictive values were calculated using magnetic resonance as the reference test. A total of 120 temporomandibular joints were examined. Based on our findings, the presence of articular sounds and physical signs are often insufficient to detect disk displacement. Imaging by static and dynamic high-resolution ultrasound demonstrates considerably lower sensitivity when compared with magnetic resonance. Some of the technical difficulties resulted from a limited access because of the presence of surrounding bone structures. Internal derangement of the temporomandibular joint: is there still a place for ultrasound? The present study does not support the recommendation of ultrasound as a conclusive diagnostic tool for internal derangements of the temporomandibular joint.
@@
Figures from the British Defence Dental Services reveal that serving personnel in the British Army have a persistently lower level of dental fitness than those in the Royal Navy or the Royal Air Force. No research had been undertaken to ascertain if this reflects the oral health of recruits joining each Service. This study aimed to pilot a process for collecting dental and sociodemographic data from new recruits to each Service and examine the null hypothesis that no differences in dental health existed. Diagnostic criteria were developed, a sample size calculated and data collected at the initial training establishments of each Service. Data for 432 participants were entered into the analysis. Recruits in the Army sample had a significantly greater prevalence of dental decay and greater treatment resource need than either of the other two Services. Army recruits had a mean number of 2.59 (2.08, 3.09) decayed teeth per recruit, compared to 1.93 (1.49, 2.39 p<0.01) in Royal Navy recruits and 1.26 (0.98, 1.53 p<0.001) in Royal Air Force recruits. Among Army recruits 62.7% were from the two most deprived quintiles of the Index of Multiple Deprivation compared to 42.5% of Royal Naval recruits and 36.6% of Royal Air Force recruits. Is there a differential in the dental health of new recruits to the British Armed Forces? A significant difference in dental health between recruits to each Service does exist and is a likely to be a reflection of the sociodemographic background from which they are drawn.
@@
Appendix F Appendix - Non-uniform transport costs Systems
We study how the uniform transport cost affects ULD Loss by comparing distillation results with and without it, as shown in Tab. 16. To emphasize the computational cost of the Wasserstein distance with non-uniform transport costs, which scales as $\mathcal{O}(n^{3}\log n)$ with $n$ being the size of the larger support, we use teacher and student models with both small and large vocabulary sizes. For small vocabularies, we select LLaMA ($32,000$ tokens) as the teacher and Pythia-410 ($50,304$ tokens) as the student. For large vocabularies, we pair LLaMA with Bloomz-560m ($250,880$ tokens). We use $5k$ samples of the SQuAD dataset due to its short answers, averaging $5.52$ tokens for LLaMA, $3.48$ for Pythia, and $3.37$ for Bloomz. For non-uniform transport costs, the Levenshtein distance and the L2 norm between embedding tokens are used to define the cost of transporting mass between points in the transport plan.
The Levenshtein distance (Lcvenshtcin, 1966) measures how different two strings or tokens are. This distance is the smallest number of operations needed to turn one string into another. These operations are: inserting a character, deleting a character, or substituting a character, each increasing the cost of transformation by $1$. For example, converting “cat” to “cats” needs one insertion, giving a distance of 1.
Before distillation training, we calculated the distance between all teacher and student tokens, resulting in a matrix $C$ of size $(|\Omega^{t}|,|\Omega^{s}|)$ with $|\Omega^{t}|$ the teacher vocabulary size and $|\Omega^{s}|)$ the student vocabulary size. Each element $C_{ij}=C_{ji}$ represents the transport cost between points $i$ and $j$ in the transport plan. This matrix was used as a cost dictionary during training in the Wasserstein distance computation Eq. 3.
Embedding token (Bojanowski et al., 2017) is a process that maps tokens (such as words or subwords) into continuous vector representations in a high-dimensional space. In this space, tokens with similar meanings or contexts are positioned closer together, reflecting their semantic similarity. We use the L2 norm between the teacher and student token embeddings to compute the cost matrix. However, since the teacher and student models do not share a common embedding space, we employ the widely used FastText subword embeddings  [^9] https://fasttext.cc/docs/en/english-vectors.html as a shared space. To achieve this, each token from the teacher and student models is mapped to its nearest token in the FastText vocabulary based on the Levenshtein distance. The corresponding FastText embedding is then assigned to the respective teacher or student token. This approach allows us to compute a matrix of L2 norm distances between all teacher and student token embedding representations. This distance matrix is then used as a cost dictionary during training in the Wasserstein distance computation.
To compute the Wasserstein distance during training, we used the POT library (Flamary et al., 2021). Extensive GPU memory management including, Offloading Optimizer/Gradient state states to CPU, activation offloading, and gradient accumulation, were employed to address the high memory demands of calculating transport plans and cost matrices. We highlight, that these experiments were conducted on A100 GPUs with 80GB of memory, using SQuAD dataset with an average label length of $4.12$ tokens. Longer sentences were infeasible due to severe memory constraints.
Table 16: Distillation using ULD Loss for LLama teacher to a Pythia-410m and Bloomz-560m student with and without uniform transport cost.
Student Transport Cost F1 score Training time
Pythia Uniform 69.56 64 minutes (×$\times 1$)
Pythia Levenshtein distance 69.32 172 minutes (×$\times 2.68$)
Pythia Embedding L2 norm 69.61 203 minutes (×$\times 3.17$)
Bloomz Uniform 72.11 59 minutes (×$\times 1$)
Bloomz Levenshtein distance 72.08 221 minutes (×$\times 3.74$)
Bloomz Embedding L2 norm 72.37 265 minutes (×$\times 4.49$)
For all experiments, in our limited setting, we do not observe a consistent improvement between models trained with ULD Loss and uniform transport costs defined by the Levenshtein distance or Embedding L2 Norm. These methods were however slower, requiring up to 4.49×4.49$\times$4.49 × the time of uniform cost training, particularly as the model vocabulary size increased, such as when distilling Llama into Bloomz. Furthermore, training with non-uniform transport costs introduced additional challenges, including pre-training steps and intensive memory management to avoid GPU “out of memory” errors. These findings emphasize the practicality of using uniform costs for maintaining training efficiency while achieving comparable performance. We encourage future work to explore novel cost matrices that balance computational complexity with potential gains in model performance.
Appendix G Appendix - Training Information